### v3

# Trace Reconstruction of First-Order Reed-Muller Codewords Using Run Statistics

Abstract
In this paper, we derive an expression for the expected number of runs in a trace of a binary sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ obtained by passing $x$ through a deletion channel that independently deletes each bit with probability $q$. We use this expression to show that if $x$ is a codeword of a first-order Reed-Muller code, and the deletion probability $q$ is $1/2$, then $x$ can be reconstructed, with high probability, from $\widetilde{O}(n^{2})$ many of its traces.
I Introduction
In the trace reconstruction problem, a sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$ is to be reconstructed exactly from multiple independent traces obtained by repeatedly passing $x$ through an i.i.d. deletion channel that independently deletes each bit with probability $q$. The goal is to determine the trace complexity of reconstruction, i.e., the minimum number of traces required to reconstruct $x$ with high probability (w.h.p.) and also to find an efficient reconstruction algorithm. The problem has its origins in the work of Levenshtein [Lev97, Lev01], but in the form described above, it was first introduced by Batu et al. [BKKM04].
[…]
Our approach is in the spirit of that of Chase [Chase20], which uses a statistic based on the appearances of a specific subsequence in the observed traces to distinguish between a pair of sequences $x$ and $y$. We use the simpler notion of runs (i.e., contiguous subsequences of like symbols) to craft our statistic. Specifically, we demonstrate that the number of runs of $\mathsf{0}$s and $\mathsf{1}$s in the observed traces can be used to differentiate between the codewords of $\mathrm{RM}(m,1)$. Doing so, we obtain an $\widetilde{O}(n^{2})$ upper bound on the (worst-case) trace complexity of sequences from $\mathrm{RM}(m,1)$ for a deletion probability of $q=\frac{1}{2}$. While runs of $\mathsf{0}$s and $\mathsf{1}$s have previously been used in reconstruction algorithms [BKKM04], [CGMR20], [KMMP21], they were invariably used as an aid in aligning the bits of a trace to the bits of the original sequence $x$. In contrast, our use of the number of runs in a trace is purely statistical.
This paper has two main contributions. The first is Theorem 1, presented in Section III, which gives an expression for the expected number of runs in a (random) trace of a given sequence $x\in\{\mathsf{0},\mathsf{1}\}^{n}$. The other is Theorem 2 (Section IV), which states that the trace complexity of reconstructing w.h.p. an arbitrary codeword from $\mathrm{RM}(m,1)$ is $\widetilde{O}(n^{2})$ when the deletion probability $q$ is $\frac{1}{2}$.
II Notation and Preliminaries
We use $[n]$ to denote the set $\{1,2,\ldots,n\}$. Binary sequences, i.e., sequences over the alphabet $\{\mathsf{0},\mathsf{1}\}$, will be denoted by $x,y,z,\omega$ etc., with $x_{i},y_{i},z_{i},\omega_{i}$ etc. denoting the $i$-th bits in the respective sequences. If $x=(x_{1},\ldots,x_{n})$, then $n$ is called the length of the sequence $x$. The concatenation $x\|y$ of two sequences $x=(x_{1},\ldots,x_{m})$ and $y=(y_{1},\ldots,y_{n})$ is the sequence $(x_{1},\ldots,x_{m},y_{1},\ldots,y_{n})$. The complement of a sequence $x$ is the sequence $\overline{x}$ having $i$-th bit $\overline{x_{i}}$, where $\overline{x_{i}}$ is the complement of the bit $x_{i}$, defined via $\overline{\mathsf{0}}=\mathsf{1}$ and $\overline{\mathsf{1}}=\mathsf{0}$. We will use the notation $\widehat{x}=x\|x$ and $\widecheck{x}=x\|\overline{x}$.
Let $x$ be a binary sequence of length $n$. The support of $x$ is the set $\mathcal{S}_{x}=\{i:x_{i}=\mathsf{1}\}$. We will use $\overline{\mathcal{S}}_{x}$ to denote the complementary set $n]\setminus\mathcal{S}_{x}=\{i:x_{i}=\mathsf{0}\}$. The Hamming weight of $x$ is $w_{H}(x)=|\mathcal{S}_{x}|$. A run of $\mathsf{0}$s of length $l$ in $x$ is a subsequence of the form $(x_{i},x_{i+1},\cdots,x_{i+l-1})$, $1\leq i\leq n-l+1$, with $x_{i-1}=\mathsf{1},x_{i}=x_{i+1}=\cdots=x_{i+l-1}=\mathsf{0},x_{i+l}=\mathsf{1}$, where to handle edge cases, we set $x_{0}=x_{n+1}=\mathsf{1}$. A run of $\mathsf{1}$s is defined analogously.
Given a binary sequence $x$, a trace of $x$ is a (random) subsequence, $T_{x}$, obtained by passing $x$ though an i.i.d. deletion channel that independently deletes each bit in $x$ with probability $q\in(0,1)$. Let $R_{x,\mathsf{0}}$ and $R_{x,\mathsf{1}}$ be the (random) number of runs of $\mathsf{0}$s and $\mathsf{1}$s, respectively, in $T_{x}$. The total number of runs in $T_{x}$ is denoted by $R_{x}=R_{x,\mathsf{0}}+R_{x,\mathsf{1}}$. Multiple traces of $x$ are denoted by $T_{x}^{1},T_{x}^{2},\ldots$, and we use $R_{x,\mathsf{0}}^{i}$, $R_{x,\mathsf{1}}^{i}$ and $R_{x}^{i}$ to denote the number of runs of $\mathsf{0}$s, number of runs of $\mathsf{1}$s, and total number of runs, respectively, in $T_{x}^{i}$.
II-A Binary First-Order Reed-Muller Codes
In this section, we record some standard facts about binary first-order Reed-Muller (RM) codes that will be used later in the paper; for more information on RM codes, see, e.g., [ASY21]. Let $\mathbb{F}_{2}$ denote the binary field. For a fixed integer $m\geq 1$, let $\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}$ be $m$ Boolean variables. For a polynomial $f\in\mathbb{F}_{2}[\mathsf{x}_{1},\mathsf{x}_{2},\ldots,\mathsf{x}_{m}]$, and a binary $m$-tuple $z=(z_{1},z_{2},\ldots,z_{m})\in\mathbb{F}_{2}^{m}$, the evaluation of $f$ at $z$ is $\text{Eval}_{z}(f)=f(z_{1},z_{2},\cdots,z_{m})$, all arithmetic taking place in $\mathbb{F}_{2}$. Then, we define the evaluation vector $\text{Eval}(f):=(\text{Eval}_{z}(f):z\in\mathbb{F}_{2}^{m})$, with the evaluation points $z\in\mathbb{F}_{2}^{m}$ appearing in the standard lexicographic order. The binary first-order RM code $\mathrm{RM}(m,1)$ is the set of all evaluation vectors $\text{Eval}(f)$ of polynomials $f$ of the form $u_{0}+u_{1}\mathsf{x}_{1}+u_{2}\mathsf{x}_{2}+\cdots+u_{m}\mathsf{x}_{m}$, with $u_{i}\in\mathbb{F}_{2}$ for all $i$. The code has blocklength $n=2^{m}$, dimension $m+1$, and minimum distance $n/2=2^{m-1}$. In fact, the code consists of the all-zeros and all-ones vectors, n1n-1 vectors of Hamming weight $n/2$ with first coordinate $\mathsf{0}$, and n1n-1 vectors of Hamming weight $n/2$ with first coordinate $\mathsf{1}$. In particular, since the all-ones vector is a codeword, $x\in\mathrm{RM}(m,1)$ implies that $\overline{x}\in\mathrm{RM}(m,1)$.
The code $\mathrm{RM}(m,1)$ can be constructed recursively from $\mathrm{RM}(m-1,1)$. Recalling our notation $\widehat{x}=x\|x$ and $\widecheck{x}=x\|\overline{x}$ from the previous subsection, we have $\mathrm{RM}(m,1)=\{\widehat{x}:x\in\mathrm{RM}(m-1,1)\}\cup\{\widecheck{x}:x\in\mathrm{RM}(m-1,1)\}$. The recursion is initialized by $\mathrm{RM}(1,1)=\mathbb{F}_{2}^{2}=\{\mathsf{0}\mathsf{0},\mathsf{0}\mathsf{1},\mathsf{1}\mathsf{0},\mathsf{1}\mathsf{1}\}$.
From the recursive construction, it follows that the first four coordinates of every codeword in $\mathrm{RM}(m,1)$, $m\geq 2$, forms a codeword in $\mathrm{RM}(2,1)$, and hence, is of the form $x\|x$ or $x\|\bar{x}$ for some $x\in\{\mathsf{0}\mathsf{0},\mathsf{0}\mathsf{1},\mathsf{1}\mathsf{0},\mathsf{1}\mathsf{1}\}$.
II-B Chernoff Bounds
In this paper, we will make much use of the following standard version of the Chernoff bounds: Let $X=\sum\limits_{i=1}^{n}X_{i}$ where $X_{i}$ is a Bernoulli random variable with $\Pr(X_{i}=\mathsf{1})=p_{i}$. Let $\mu=\mathbb{E}[X]=\sum_{i=1}^{n}p_{i}$. We then have
  • for $\delta>0$, $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/(2+\delta)}$; in particular, for $0<\delta<1$, $\Pr(X\geq(1+\delta)\mu)\leq e^{-\mu\delta^{2}/3}$
  • for $0<\delta<1,\ \Pr(X\leq(1-\delta)\mu)\leq e^{-\mu\delta^{2}/2}$.
II-C Hoeffding’s Inequality
Let $X_{1},\cdots,X_{n}$ be independent random variables such that $a_{i}\leq X_{i}\leq b_{i}$ a.s. If $X=\sum\limits_{i=1}^{n}X_{i}$, then Hoeffding’s inequality states that
$\Pr(|X-\mathbb{E}[X]|\geq t)\leq 2\exp\biggl(-\frac{2t^{2}}{\sum\limits_{i=1}^{n}(b_{i}-a_{i})^{2}}\biggr)$
III Expected Number of Runs in a Trace
One of the main contributions of this paper is an explicit expression for the expected number of runs of $\mathsf{0}$s and $\mathsf{1}$s in a trace of a given binary sequence $x$ of length $n$.
Theorem 1.
Given a binary sequence $x$, the expected numbers of runs of $\mathsf{0}$s and $\mathsf{1}$s in a trace $T_{x}$ obtained by independently deleting each bit in $x$ with probability $q$ are given by
$\mathbb{E}[R_{x,\mathsf{0}}]=(1-q)\cdot\left(n-w_{H}(x)-\frac{1-q}{q}\sum\limits_{i,j\in\overline{\mathcal{S}}_{x}:i<j}q^{j-i}\right)$
and
$\mathbb{E}[R_{x,\mathsf{1}}]=(1-q)\cdot\left(w_{H}(x)-\frac{1-q}{q}\sum\limits_{i,j\in{\mathcal{S}}_{x}:i<j}q^{j-i}\right),$
respectively. Consequently, $\mathbb{E}[R_{x}]=\mathbb{E}[R_{x,\mathsf{0}}]+\mathbb{E}[R_{x,\mathsf{1}}]$ is given by
$(1-q)\cdot\left(n-\frac{1-q}{q}\left(\sum\limits_{i,j\in\overline{\mathcal{S}}_{x}:i<j}q^{j-i}+\sum\limits_{i,j\in{\mathcal{S}}_{x}:i<j}q^{j-i}\right)\right).$
Proof.
When $x$ is passed through an i.i.d. deletion channel with deletion probability $q$, the action of the channel is captured by a (random) binary sequence $\omega$ of length $n$, in which $\omega_{j}=\mathsf{1}$ iff the channel deletes $x_{j}$, and $\Pr(\omega)=q^{w_{H}(\omega)}(1-q)^{n-w_{H}(\omega)}$. The resulting trace is $T_{x}(\omega)={(x_{j})}_{j\in\overline{\mathcal{S}}_{\omega}}$.
We only prove the result for $R_{x,\mathsf{0}}$, since the result for $R_{x,\mathsf{1}}$ is obtained entirely analogously by exchanging the roles of $\mathsf{0}$s and $\mathsf{1}$s. So, let us consider the runs of $\mathsf{0}$s in $T_{x}(\omega)$. Observe that the $j$-th bit in the sequence $x$ will start a new run of $\mathsf{0}$s in $T_{x}(\omega)$ iff $x_{j}=0$, $\omega_{j}=\mathsf{0}$ (so that $x_{j}$ is not deleted), and the previous undeleted bit in $x$ (if it exists) is a $\mathsf{1}$. We then define $n$ indicator random variables $I_{1},I_{2},\ldots,I_{n}$ such that $I_{j}=1$ iff the $j$-th bit in $x$ starts a new run of $\mathsf{0}$s in the random trace $T_{x}$. With this, we have $R_{x,\mathsf{0}}=\sum\limits_{j=1}^{n}I_{j}$. In fact, $R_{x,\mathsf{0}}=\sum_{j\in\overline{S}_{x}}I_{j}$, since $I_{j}=0$ if $x_{j}=\mathsf{1}$. Thus, $\mathbb{E}[R_{x,\mathsf{0}}]=\sum_{j\in\overline{S}_{x}}\Pr(I_{j}=1)$. We next determine $\Pr(I_{j}=1)$ for $j\in\overline{\mathcal{S}}_{x}$.
When $j\in\overline{\mathcal{S}}_{x}$, we have
$\Pr(I_{j}=1)$ $=\Pr(I_{j}=1,\omega_{j}=0)$
$=(1-q)\cdot\Pr(I_{j}=1\mid\omega_{j}=0)$
$=\ (1-q)\cdot\bigl(1-\Pr(I_{j}=0\mid\omega_{j}=0)\bigr).$
Now, given that $x_{j}=\mathsf{0}$ is undeleted, we can have $I_{j}=0$ iff the previous undeleted bit in $x$ is a $\mathsf{0}$. In other words, conditioned on $x_{j}=0$ being undeleted, we have $I_{j}=0$ iff the event $\mathcal{E}_{i}:=\{\omega_{i}=\mathsf{0},\text{ but }\omega_{i+1}=\cdots=\omega_{j-1}=\mathsf{1}\}$ occurs for some $i<j$ such that $x_{i}=\mathsf{0}$. Since the events $\mathcal{E}_{i}$ are mutually disjoint and the $\omega_{\ell}$’s are i.i.d., we obtain that for $j\in\overline{\mathcal{S}}_{x}$,
$\Pr(I_{j}=0\mid\omega_{j}=0)$ $=\sum_{i\in\overline{\mathcal{S}}_{x}:i<j}\Pr(\mathcal{E}_{i})$
$=\sum_{i\in\overline{\mathcal{S}}_{x}:i<j}(1-q)\,q^{j-i-1}.$
Putting it all together, we obtain
$\mathbb{E}[R_{x,\mathsf{0}}]$ $=(1-q)\cdot\sum_{j\in\overline{S}_{x}}\bigl(1-\Pr(I_{j}=0\mid\omega_{j}=0)\bigr)$
$=(1-q)\cdot\left(n-w_{H}(x)-\frac{1-q}{q}\sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}q^{j-i}\right)$
as desired. ∎
For the special case when $q=\frac{1}{2}$, we have
$\mathbb{E}[R_{x}]=\frac{1}{2}\left(n-\sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}{\biggl(\frac{1}{2}\biggr)}^{j-i}-\sum_{i,j\in{\mathcal{S}}_{x}:i<j}{\biggl(\frac{1}{2}\biggr)}^{j-i}\right)\,.$ (1)
IV Trace Reconstruction for $\mathrm{RM}(m,1)$
From here on, we consider only the case of deletion probability $q=\frac{1}{2}$. Our main result is that, using the average number of runs in a trace as a statistic, one can reconstruct, with high probability (w.h.p.), any $x\in\mathrm{RM}(m,1)$ using only $\widetilde{O}(n^{2})$ of its traces, where $n=2^{m}$ is the blocklength of the code.
Theorem 2.
For deletion probability $q=\frac{1}{2}$, any codeword $x\in\mathrm{RM}(m,1)$ can be correctly reconstructed from $\widetilde{O}(n^{2})$ traces, with probability $\geq 1-1/\mathrm{poly}(n)$.
We first describe the proof strategy. Recall that $\mathrm{RM}(m,1)$ consists of $2n=2^{m+1}$ codewords, of which half begin with a $\mathsf{0}$ and the other half begin with a $\mathsf{1}$. By examining the first bit of $O(\log n)$ traces, the first bit of $x$ can be correctly determined w.h.p. — see Claim 1 below. It then remains to distinguish between the $n$ codewords of having $x_{1}$ as the first bit. This, as we will show, can be achieved w.h.p. using run statistics obtained from $O(n^{2}\log n)$ traces. What makes this possible is the following property of the $\mathrm{RM}(m,1)$ code, which may be of independent interest.
Proposition 1.
If $x$ and $y$ are two distinct codewords in $\mathrm{RM}(m,1)$, $m\geq 4$, that agree in their first bit (i.e., $x_{1}=y_{1}$), then $\bigl|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]\bigr|\geq 0.028$.
The proof of Proposition 1 is by induction on $m$, starting with the base case of $m=4$. We defer the proof to Section IV-B. We first give a proof of Theorem 2, assuming the validity of Proposition 1.
IV-A Proof of Theorem 2
The algorithm used for reconstructing $x\in\mathrm{RM}(m,1)$ from $k$ non-empty traces $T_{x}^{1},T_{x}^{2},\ldots,T_{x}^{k}$ is described as follows:
Step 1:
Consider the first $\ell<k$ traces. If more than $\ell/2$ of these traces start with a $\mathsf{1}$, set $\mathsf{b}=\mathsf{1}$; else, set $\mathsf{b}=\mathsf{0}$.
Step 2:
Compute $\overline{R}:=\frac{1}{k}\sum_{i=1}^{k}(\#\text{ runs in trace }T_{x}^{i})$. Find a $z\in\mathrm{RM}(m,1)$ with first bit $z_{1}=\mathsf{b}$ (as obtained in Step 1) that minimizes $|\mathbb{E}[R_{z}]-\overline{R}|$ among all such $z$. Output $z$ and stop.
We will prove that if the algorithm takes $k=O(n^{2}\log n)$ random traces as input, and we set $\ell=O(\log n)$ in Step 1, then with probability at least $1-1/\mathrm{poly}(n)$, the codeword output by the algorithm is equal to the $x$ from which the $k$ traces were obtained. In what follows, $\mathbb{P}_{x}$ denotes the probability law of the traces $T_{x}^{1},T_{x}^{2},\ldots,T_{x}^{k}$, and $Z$ is the random variable representing the output of the reconstruction algorithm. Reconstruction Error is the event $Z\neq x$, and $\mathcal{E}_{j}$, for $j=1,2$, is the event “Error occurs in Step jj”. Note that
$\mathbb{P}_{x}(\textsc{Reconstruction Error})\leq\mathbb{P}_{x}(\mathcal{E}_{1})+\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})$ (2)
The terms on the right-hand side are bounded in Claims 1 and 2 below.
Claim 1.
If $\ell=O(\log n)$ traces are used in Step 1, then for any $x\in\mathrm{RM}(m,1)$, we have $\mathbb{P}_{x}(\mathcal{E}_{1})\leq n^{-c_{1}}$ for some constant $c_{1}>0$.
Proof.
The number of traces among $T_{x}^{1},T_{x}^{2},\ldots,T_{x}^{\ell}$ having first bit $\mathsf{1}$ is $A=\sum_{i=1}^{\ell}A_{i}$ where
$A_{i}=\begin{cases}1&\text{if the first bit in $T_{x}^{i}$ is }\mathsf{1},\\ 0&\text{otherwise.}\end{cases}$
Suppose that the first bit of $x\in\mathrm{RM}(m,1)$ is $\mathsf{1}$. If the second bit is also $\mathsf{1}$, then $\mathbb{E}[A_{i}]\geq 1-\mathbb{P}_{x}($the first two bits of $x$ are deleted)=3/4)=3/4. If the second bit of $x$ is $\mathsf{0}$, then the next two bits are either $\mathsf{1}\mathsf{0}$ or $\mathsf{0}\mathsf{1}$ (see Section II-A). In this case, $\mathbb{E}[A_{i}]\geq\mathbb{P}_{x}($one of the two $\mathsf{1}$s in the first four bits of $x$ is the first bit of trace $T_{x}^{i})\geq(1/2+1/16)=9/16$. Thus, when the first bit of $x$ is $\mathsf{1}$, then $\mathbb{E}[A]\geq\frac{9}{16}\ell$. Via the Chernoff bound, we obtain $\mathbb{P}_{x}(\mathcal{E}_{1})=\mathbb{P}_{x}(A\leq\frac{1}{2}\ell)\leq e^{-\frac{\ell}{288}}$.
Next, suppose that the first bit of $x\in\mathrm{RM}(m,1)$ is $\mathsf{0}$. Note that the number of traces having first bit $\mathsf{0}$ is $\ell-A$. So, exchanging the roles of $\mathsf{0}$ and $\mathsf{1}$ in the above argument, we obtain $\mathbb{P}_{x}(\ell-A\leq\frac{1}{2}\ell)\leq e^{-\frac{\ell}{288}}$. Then, $\mathbb{P}_{x}(\mathcal{E}_{1})=\mathbb{P}_{x}(A>\frac{1}{2}\ell)\leq\mathbb{P}_{x}(A\geq\frac{1}{2}\ell)\leq e^{-\frac{\ell}{288}}$.
Thus, for any $x\in\mathrm{RM}(m,1)$, we have $\mathbb{P}_{x}(\mathcal{E}_{1})\leq e^{-\frac{\ell}{288}}$. Hence, if $\ell=c_{0}\log n$, then $\mathbb{P}_{x}(\mathcal{E}_{1})\leq n^{-c_{0}/288}$. ∎
Claim 2.
If $k=O(n^{2}\log n)$ traces are given as input to the reconstruction algorithm, then for any $x\in\mathrm{RM}(m,1)$, we have $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq 2\cdot n^{-c_{2}}$ for some constant $c_{2}>0$.
Proof.
Since $\mathbb{P}_{x}(\mathcal{E}_{2}\cap\mathcal{E}_{1}^{c})\leq\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})$, it suffices to bound the latter probability. Let $\mathbb{E}[R_{x}]=r_{x}\leq n$. From Proposition 1, for any other $y\in\mathrm{RM}(m,1)$ having the same first bit as $x$, we have $|r_{x}-\mathbb{E}[R_{y}]|\geq\Delta:=0.028$. So, if the reconstruction algorithm does not make an error in Step 1 (i.e., given $\mathcal{E}_{1}^{c}$), then at the end of Step 2, it will output $Z=x$ if $r_{x}-\Delta/2<\overline{R}<r_{x}+\Delta/2$. Consequently, given $\mathcal{E}_{1}^{c}$, if error event $\mathcal{E}_{2}$ occurs, then $|\overline{R}-r_{x}|\geq\Delta/2$.
Note that $k\overline{R}=\sum_{i=1}^{k}R_{x}^{i}$. Since $0\leq R_{x}^{i}\leq n$, using Hoeffding’s inequality, we obtain
$\Pr(|\overline{R}-r_{x}|\geq\Delta/2)=\Pr(|k\overline{R}-kr_{x}|\geq k\Delta/2)\leq 2\exp\biggl(-\frac{k\Delta^{2}}{2n^{2}}\biggr)$
Taking $k=c_{3}n^{2}\log n$ (for suitable $c_{3}>0$), we obtain $\mathbb{P}_{x}(\mathcal{E}_{2}\mid\mathcal{E}_{1}^{c})\leq 2\cdot e^{-\frac{c_{3}\Delta^{2}}{2}\log n}=2n^{-c_{2}}$. ∎
The proof of Theorem 2 is completed by bounding the probability of Reconstruction Error in (2) using the bounds in Claims 1 and 2.
IV-B Proof of Proposition 1
Recall from Eq. (1) that, for a sequence $x$, $\mathbb{E}[R_{x}]=\frac{1}{2}(n-\alpha_{x})$, where we define
$\alpha_{x}=\sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{j-i}+\sum_{i,j\in{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{j-i}.$
For the purposes of this proof, we require a few more coefficients to be defined:
$\beta_{x}\$$=\ \sum_{i,j\in\overline{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}+\sum_{i,j\in{\mathcal{S}}_{x}:i<j}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}$
$\gamma_{x}\$$=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{n-(j-i)}$
$\delta_{x}\$$=\ \sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{j-i}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}\biggl(\frac{1}{2}\biggr)^{j-i}.$
Our proof relies on the technical lemma stated next. While the coefficient $\delta_{x}$ does not appear in the statement of the lemma, it will be needed in its proof.
Lemma 3.
For $m\geq 4$ and $n=2^{m}$, conditions (C1)(C3) below hold for any pair of distinct codewords $x,y\in\mathrm{RM}(m,1)$ having the same first bit, and condition (C4) holds for any $x\in\mathrm{RM}(m,1)$:
  • (C1) $|\alpha_{x}-\alpha_{y}|\ \geq\ 0.06-3\sum\limits_{k=16}^{n/2}k\,(\frac{1}{2})^{k}$
  • (C2) $|\alpha_{x}-\alpha_{y}|\,-\,|\beta_{x}-\beta_{y}|\ \geq\ 0$
  • (C3) $|\alpha_{x}-\alpha_{y}|\,-\,|\beta_{x}-\gamma_{y}|\ \geq\ 0.06-4\sum\limits_{k=16}^{n/2}k\,(\frac{1}{2})^{k}$
  • (C4) $|\beta_{x}-\gamma_{x}|\ \geq\ 0.06-3\sum\limits_{k=16}^{n/2}k\,(\frac{1}{2})^{k}$
We remark here that we only require condition (C1) of the lemma to complete the proof of Proposition 1. The remaining conditions (C2)–(C4) are auxiliary conditions needed in our induction-based proof of the lemma, which we provide in the next section.
From (C1), we obtain that for any pair of distinct codewords $x,y\in\mathrm{RM}(m,1)$, $m\geq 4$, having the same first bit,
$|\alpha_{x}-\alpha_{y}|\geq 0.057,$ (3)
since $\sum\limits_{k=16}^{n/2}k\,(\frac{1}{2})^{k}\leq\sum\limits_{k=16}^{\infty}k\,(\frac{1}{2})^{k}=17\cdot(\frac{1}{2})^{15}<0.001$. Now, let $x$ and $y$ be two distinct codewords of $\mathrm{RM}(m,1)$, $m\geq 4$, such that $x_{1}=y_{1}$. Via (3), we obtain
$\bigl|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]\bigr|\ =\ \frac{1}{2}|\alpha_{x}-\alpha_{y}|\ \geq\ \frac{1}{2}\cdot 0.057\ =\ 0.0285,$
thus establishing Proposition 1.
V Proof of Lemma 3
The proof is by induction on $m$, with $m=4$ as the base case. We first argue that it suffices to prove the lemma for $x,y\in\mathrm{RM}(m,1)$ with $x_{1}=y_{1}=0$. This is because $\overline{\mathcal{S}}_{x}=\mathcal{S}_{\overline{x}}$, from which it is easy to check that $\alpha_{x}=\alpha_{\overline{x}}$, $\beta_{x}=\beta_{\overline{x}}$ and $\gamma_{x}=\gamma_{\overline{x}}$. Therefore, if codewords $x,y\in\mathrm{RM}(m,1)$ satisfy conditions (C1)–(C4), then so do the codewords $\overline{x},\overline{y}$.
TABLE I: Values of $\alpha_{x},\beta_{x},\gamma_{x}$ for $x\in\mathrm{RM}(4,1)$ with $x_{1}=0$
Codeword $x$ $\alpha_{x}$ $\beta_{x}$ $\gamma_{x}$
$00000000$ $\frac{458753}{32768}$ $\frac{65519}{32768}$ 0
0000000011111111 $\frac{769}{64}$ $\frac{247}{16384}$ $\frac{65025}{32768}$
$00001111$ $\frac{8929}{1024}$ $\frac{1811}{8192}$ $\frac{58275}{32768}$
0000111111110000 $\frac{336353}{32768}$ $\frac{57869}{32768}$ $\frac{3825}{16384}$
$00110011$ $\frac{23593}{4096}$ $\frac{655}{1024}$ $\frac{44559}{32768}$
0011001111001100 $\frac{212153}{32768}$ $\frac{44369}{32768}$ $\frac{10575}{16384}$
$00111100$ $\frac{251033}{32768}$ $\frac{41939}{32768}$ $\frac{5895}{8192}$
0011110011000011 $\frac{29101}{4096}$ $\frac{11857}{16384}$ $\frac{41805}{32768}$
$01010101$ $\frac{36409}{8192}$ $\frac{7279}{8192}$ $\frac{36403}{32768}$
0101010110101010 $\frac{152861}{32768}$ $\frac{36341}{32768}$ $\frac{14589}{16384}$
$01011010$ $\frac{164861}{32768}$ $\frac{35591}{32768}$ $\frac{3741}{4096}$
0101101010100101 $\frac{39809}{8192}$ $\frac{14983}{16384}$ $\frac{35553}{32768}$
$01100110$ $\frac{175637}{32768}$ $\frac{34067}{32768}$ $\frac{7863}{8192}$
0110011010011001 $\frac{43259}{8192}$ $\frac{15733}{16384}$ $\frac{34053}{32768}$
$01101001$ $\frac{42179}{8192}$ $\frac{3967}{4096}$ $\frac{33783}{32768}$
0110100110010110 $\frac{170741}{32768}$ $\frac{33761}{32768}$ $\frac{15879}{16384}$
Table I lists the $\alpha_{x}$, $\beta_{x}$ and $\gamma_{x}$ coefficients needed to verify the lemma in the base case of $\mathrm{RM}(4,1)$. Indeed, from the table it can be verified that for any pair of distinct $x,y\in\mathrm{RM}(4,1)$ with $x_{1}=y_{1}=\mathsf{0}$, we have
$|\alpha_{x}-\alpha_{y}|\geq\frac{2025}{32768}$ $\geq 0.0617,$
$|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\beta_{y}|$ $\geq 0,$
$|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\gamma_{y}|\geq\frac{2003}{32768}$ $\geq 0.0611,$
$|\beta_{x}-\gamma_{x}|\geq\frac{2003}{32768}$ $\geq 0.0611.$
Thus, the lemma holds for $\mathrm{RM}(4,1)$.
Now, assuming that the lemma holds for $\mathrm{RM}(m,1)$ for some $m\geq 4$, we will show that it holds for $\mathrm{RM}(m+1,1)$ as well. Recall that codewords of $\mathrm{RM}(m+1,1)$ are of the form $\widehat{x}=x\|x$ or $\widecheck{x}=x\|\overline{x}$ for some $x\in\mathrm{RM}(m,1)$. The coefficients $\alpha,\beta,\gamma$ for $\widehat{x}$ and $\widecheck{x}$ can be expressed in terms of the coefficients for $x$, as stated in the lemma below.
Lemma 4.
For a binary sequence $x$ of length $n$, we have
  1. (a)
    $\alpha_{\widehat{x}}\ =\ 2\alpha_{x}+\beta_{x}+{(\frac{1}{2})}^{n}\alpha_{x}+n{(\frac{1}{2})}^{n}$
  2. (b)
    $\alpha_{\widecheck{x}}\ =\ 2\alpha_{x}+\gamma_{x}+{(\frac{1}{2})}^{n}\delta_{x}$
  3. (c)
    $\beta_{\widehat{x}}\ =\beta_{x}+{(\frac{1}{2})}^{n-1}\beta_{x}+{(\frac{1}{2})}^{n}\alpha_{x}+n{(\frac{1}{2})}^{n}$
  4. (d)
    $\beta_{\widecheck{x}}\ =\gamma_{x}+{(\frac{1}{2})}^{n-1}\beta_{x}+{(\frac{1}{2})}^{n}\delta_{x}$
  5. (e)
    $\gamma_{\widehat{x}}\ =\gamma_{x}+{(\frac{1}{2})}^{n-1}\gamma_{x}+{(\frac{1}{2})}^{n}\delta_{x}$
  6. (f)
    $\gamma_{\widecheck{x}}\ =\beta_{x}+{(\frac{1}{2})}^{n-1}\gamma_{x}+{(\frac{1}{2})}^{n}\alpha_{x}+n{(\frac{1}{2})}^{n}$
The expressions in the lemma are obtained through straightforward manipulations. As an illustration, we prove part (b) in Appendix A.
We need to prove conditions (C1)–(C3) in Lemma 3 for each pair of distinct codewords in $\mathrm{RM}(m+1,1)$ having the same first bit, while we need to prove condition (C4) for each individual codeword in $\mathrm{RM}(m+1,1)$. For pairs of distinct codewords, there are four cases to consider:
  • Case 1: The codewords are $\widehat{x}=x\|x$ and $\widecheck{x}=x\|\overline{x}$ for some $x\in\mathrm{RM}(m,1)$
  • Case 2: The codewords are $\widehat{x}=x\|x$ and $\widehat{y}=y\|y$ for distinct $x,y\in\mathrm{RM}(m,1)$ sharing the same first bit.
  • Case 3: The codewords are $\widecheck{x}=x\|\overline{x}$ and $\widecheck{y}=y\|\overline{y}$ for distinct $x,y\in\mathrm{RM}(m,1)$ sharing the same first bit.
  • Case 4: The codewords are $\widehat{x}=x\|x$ and $\widecheck{y}=y\|\overline{y}$ for distinct $x,y\in\mathrm{RM}(m,1)$ sharing the same first bit.
In each of these cases, our proof of conditions (C1)–(C3) makes use of some properties of the $\alpha,\beta,\gamma,\delta$ coefficients, which we record here. For any binary sequence $x$ of length $n$,
$\alpha_{x}+\delta_{x}$ $=\sum_{i,j\in[n]:i<j}{\biggl(\frac{1}{2}\biggr)}^{j-i}=:\ a(n),$
$\beta_{x}+\gamma_{x}$ $=\sum_{i,j\in[n]:i<j}{\biggl(\frac{1}{2}\biggr)}^{n-(j-i)}=:\ b(n).$
In particular, the sums on the left-hand side above depend on the sequence $x$ only via its length $n$. Note further that
$a(n)$ $=\sum_{k=1}^{n-1}(n-k){\biggl(\frac{1}{2}\biggr)}^{k}=n-2+{\biggl(\frac{1}{2}\biggr)}^{n-1}\leq\ n$ (4)
$b(n)$ $=\sum_{k=1}^{n-1}k{\biggl(\frac{1}{2}\biggr)}^{k}=\ 2-(n+1){\biggl(\frac{1}{2}\biggr)}^{n-1}\ \leq\ 2$ (5)
Therefore,
$|\alpha_{x}-\delta_{x}|\$$\leq\ \alpha_{x}+\delta_{x}\ =\ a(n)\ \leq\ n$ (6)
$|\beta_{x}-\gamma_{x}|\$$\leq\ \beta_{x}+\gamma_{x}\ =\ b(n)\ \leq\ 2.$ (7)
Proof of (C1)–(C3) in Case 1.
For (C1), we argue as follows: via Lemma 4(a),(b),
$|\alpha_{\widehat{x}}-\alpha_{\widecheck{x}}|$ $=\left|\beta_{x}-\gamma_{x}+{\textstyle{(\frac{1}{2})}^{n}}(\alpha_{x}-\delta_{x})+n{\textstyle(\frac{1}{2})}^{n}\right|$
$\geq|\beta_{x}-\gamma_{x}|-{\textstyle{(\frac{1}{2})}^{n}|\alpha_{x}-\delta_{x}|-n{(\frac{1}{2})}^{n}}$
$\stackrel{{\scriptstyle\mathrm{(i)}}}{{\geq}}{\textstyle 0.06-3\sum\limits_{k=16}^{n/2}k{(\frac{1}{2})}^{k}-2n{(\frac{1}{2})}^{n}}$
$\geq{\textstyle 0.06-3\sum\limits_{k=16}^{n}k{(\frac{1}{2})}^{k}}$ (8)
where to obtain the inequality labeled (i), we used (6) and the induction hypothesis that (C4) holds for $\mathrm{RM}(m,1)$.
To show (C2), we merely observe from Lemma 4(a)–(d) that $\alpha_{\widehat{x}}-\alpha_{\widecheck{x}}=\beta_{\widehat{x}}-\beta_{\widecheck{x}}$.
For (C3), we first note that, from Lemma 4(c),(f), we have
$|\beta_{\widehat{x}}-\gamma_{\widecheck{x}}|={\textstyle{(\frac{1}{2})}^{n-1}\,|\beta_{x}-\gamma_{x}|\stackrel{{\scriptstyle\mathrm{(ii)}}}{{\leq}}{(\frac{1}{2})}^{n-1}\cdot 2\leq n{(\frac{1}{2})}^{n}},$
where the inequality labeled (ii) is by virtue of (7). Then, using (8), we get
$|\alpha_{\widehat{x}}-\alpha_{\widecheck{x}}|-|\beta_{\widehat{x}}-\gamma_{\widecheck{x}}|$ $\geq{\textstyle 0.06-3\sum\limits_{k=16}^{n}k{(\frac{1}{2})}^{k}-n{(\frac{1}{2})}^{n}}$
$\geq{\textstyle 0.06-4\sum\limits_{k=16}^{n}k{(\frac{1}{2})}^{k}},$
which is precisely (C3). ∎
The proofs of (C1)–(C3) in the remaining Cases 2–4 are more involved, and are given in Appendix B.
Finally, we prove that condition (C4) holds for any codeword in $\mathrm{RM}(m+1,1)$.
Proof of (C4).
First, consider $\widehat{x}=x\|x$ for some $x\in\mathrm{RM}(m,1)$. From Lemma 4(c),(e), we have
[…]
$={\textstyle\left|\bigl(1+{(\frac{1}{2})}^{n-1}\bigr)(\beta_{x}-\gamma_{x})+{(\frac{1}{2})}^{n}(\alpha_{x}-\delta_{x})+n{(\frac{1}{2})}^{n}\right|}$
$\geq{\textstyle|\beta_{x}\!-\!\gamma_{x}|-{(\frac{1}{2})}^{n-1}|\beta_{x}\!-\!\gamma_{x}|-{(\frac{1}{2})}^{n}|\alpha_{x}\!-\!\delta_{x}|-n{(\frac{1}{2})}^{n}}$
$\stackrel{{\scriptstyle\mathrm{(iii)}}}{{\geq}}{\textstyle 0.06-3\sum\limits_{k=16}^{n/2}k{(\frac{1}{2})}^{k}-3n{(\frac{1}{2})}^{n}}$
$\geq{\textstyle 0.06-3\sum\limits_{k=16}^{n}k{(\frac{1}{2})}^{k}}$
where to obtain the inequality labeled (iii), we used (6), (7), and the induction hypothesis that (C4) holds for $\mathrm{RM}(m,1)$.
Next, consider $\widecheck{x}=x\|\overline{x}$ for some $x\in\mathrm{RM}(m,1)$. Using similar arguments as above,
$|\beta_{\widecheck{x}}$ $-\gamma_{\widecheck{x}}|$
$={\textstyle\left|\bigl(1-{(\frac{1}{2})}^{n-1}\bigr)(\gamma_{x}-\beta_{x})-{(\frac{1}{2})}^{n}(\alpha_{x}-\delta_{x})-n{(\frac{1}{2})}^{n}\right|}$
$\geq{\textstyle|\gamma_{x}\!-\!\beta_{x}|-{(\frac{1}{2})}^{n-1}|\gamma_{x}\!-\!\beta_{x}|-{(\frac{1}{2})}^{n}|\alpha_{x}\!-\!\delta_{x}|-n{(\frac{1}{2})}^{n}}$
$\geq{\textstyle 0.06-3\sum\limits_{k=16}^{n/2}k{(\frac{1}{2})}^{k}-3n{(\frac{1}{2})}^{n}}$
$\geq{\textstyle 0.06-3\sum\limits_{k=16}^{n}k{(\frac{1}{2})}^{k}}$
Thus, (C4) holds for all codewords in $\mathrm{RM}(m+1,1)$. ∎
This completes the proof of Lemma 3.
VI Conclusion
In this paper, we derived, for a given binary sequence $x$, an expression for the expected number of runs, $\mathbb{E}[R_{x}]$, in a random trace obtained by passing $x$ through an i.i.d. deletion channel with deletion probability $q$. We then proved that, when $q=\frac{1}{2}$, the average number of runs in $\widetilde{O}(n^{2})$ traces can be used as a statistic to distinguish (w.h.p.) between the codewords of a first-order Reed-Muller code of blocklength $n$. To prove this, we used a key property of $\mathrm{RM}(m,1)$: for distinct codewords $x,y\in\mathrm{RM}(m,1)$ that share the same first bit, $|\mathbb{E}[R_{x}]-\mathbb{E}[R_{y}]|\geq\Delta$, for some constant $\Delta>0$ that does not depend on $m$. This property does not hold, for example, for higher-order RM codes $\mathrm{RM}(m,r)$, $r\geq 2$, so $\widetilde{O}(n^{2})$ traces might not be sufficient to reconstruct $x\in\mathrm{RM}(m,r)$ (up to complementation) solely using the average number of runs in the traces as a statistic. Nonetheless, it should be possible to refine this statistic further or combine it with other statistics to better effect.
Appendix A Proof of Lemma 4(b)
We start with
$\alpha_{\widecheck{x}}=\sum_{i,j\in\overline{\mathcal{S}}_{\widecheck{x}}:i<j}{\biggl(\frac{1}{2}\biggr)}^{j-i}+\sum_{i,j\in{\mathcal{S}}_{\widecheck{x}}:i<j}{\biggl(\frac{1}{2}\biggr)}^{j-i}.$
Each of these summations is over indices $1\leq i<j\leq 2n$. We split each summation into three parts:
  1. 1.
    a summation over indices $1\leq i<j\leq n$,
  2. 2.
    a summation over indices $n+1\leq i<j\leq 2n$, and
  3. 3.
    a summation over indices $1\leq i\leq n,\ n+1\leq j\leq n$.
Thus,
$\sum_{\begin{subarray}{c}i,j\in\overline{\mathcal{S}}_{\widecheck{x}}:\\ i<j\end{subarray}}$ ${\biggl(\frac{1}{2}\biggr)}^{j-i}$
$=\sum_{\begin{subarray}{c}i,j\in\overline{\mathcal{S}}_{\widecheck{x}}:\\ 1\leq i<j\leq n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{j-i}+\sum_{\begin{subarray}{c}i,j\in\overline{\mathcal{S}}_{\widecheck{x}}:\\ n+1\leq i<j\leq 2n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{j-i}$
$\quad\quad+\sum_{\begin{subarray}{c}i,j\in\overline{\mathcal{S}}_{\widecheck{x}}:\\ 1\leq i\leq n,\,n+1\leq j\leq 2n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{j-i}$
$=\sum_{\begin{subarray}{c}i,j\in\overline{\mathcal{S}}_{x}:\\ 1\leq i<j\leq n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{j-i}+\sum_{\begin{subarray}{c}i,j\in\mathcal{S}_{x}:\\ 1\leq i<j\leq n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{j-i}$
$\quad\quad+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ 1\leq i,j\leq n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}$
$=\alpha_{x}+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ 1\leq i,j\leq n\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}$
$=\alpha_{x}+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i>j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n-(i-j)}$
Analogously,
$\sum_{\begin{subarray}{c}i,j\in{\mathcal{S}}_{\widecheck{x}}:\\ i<j\end{subarray}}$ ${\biggl(\frac{1}{2}\biggr)}^{j-i}$
$=\alpha_{x}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i>j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n-(i-j)}.$
Therefore,
$\alpha_{\widecheck{x}}$ $=2\alpha_{x}+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i<j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}+\sum_{\begin{subarray}{c}i\in\overline{\mathcal{S}}_{x},j\in\mathcal{S}_{x}:\\ i>j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n-(i-j)}$
$\quad\quad+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i<j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n+j-i}+\sum_{\begin{subarray}{c}i\in{\mathcal{S}}_{x},j\in\overline{\mathcal{S}}_{x}:\\ i>j\end{subarray}}{\biggl(\frac{1}{2}\biggr)}^{n-(i-j)}$
$=2\alpha_{x}+{\biggl(\frac{1}{2}\biggr)}^{n}\delta_{x}+\gamma_{x}.$
Appendix B Proofs of (C1)–(C3) in Cases 2–4
First let us demonstrate a few more properties of coefficients $\alpha,\beta,\gamma,\delta$ that we will be using in our proofs. As can be observed from their definitions, coefficients $\alpha,\beta,\gamma,\delta$ are non-negative. Thus, for a sequence $x$ of length $n$, using (4) and (5), we get $0\leq\alpha_{x}\leq n,\ 0\leq\delta_{x}\leq n,\ 0\leq\beta_{x}\leq 2,\ 0\leq\gamma_{x}\leq 2$. As a result, for any two sequences $x$ and $y$ of length $n$,
$|\alpha_{x}-\alpha_{y}|\leq n$ (9)
$|\alpha_{x}-\delta_{y}|\leq n$ (10)
$|\beta_{x}-\beta_{y}|\leq 2$ (11)
$|\beta_{x}-\gamma_{y}|\leq 2$ (12)
Moreover,
$(\beta_{x}-\beta_{y})$ $=(b(n)-\gamma_{x}-b(n)+\gamma_{y})=-(\gamma_{x}-\gamma_{y})$ (13)
$(\alpha_{x}-\alpha_{y})$ $=(a(n)-\delta_{x}-a(n)+\delta_{y})=-(\delta_{x}-\delta_{y})$ (14)
$|\beta_{x}-\gamma_{y}|$ $=|b(n)-\gamma_{x}-b(n)+\beta_{y}|=|\beta_{y}-\gamma_{x}|$ (15)
B-A Proof of (C1)–(C3) in Case 2
Using Lemma 4(a),(c), we have
[…]
To show (C1) and (C2), we will split this case into two subcases and verify (C1) and (C2) for each.
  1. (i)
    If $(\alpha_{x}-\alpha_{y})$ and $(\beta_{x}-\beta_{y})$ have the same sign, then
[…]
    For (C2),
[…]
    $\stackrel{{\scriptstyle\mathrm{(iv)}}}{{\leq}}\textstyle{\bigl(1+3(\frac{1}{2})^{n}\bigr)|\alpha_{x}-\alpha_{y}|}$
[…]
    where we derive inequality labeled (iv) by applying the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$, while inequality labeled (v) is obtained directly from (16).
  2. (ii)
    If $(\alpha_{x}-\alpha_{y})$ and $(\beta_{x}-\beta_{y})$ have different signs, then
[…]
    where we get the above inequality using the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$.
    For (C2),
[…]
    $=\textstyle{|\beta_{x}\!-\beta_{y}|+(\frac{1}{2})^{n}\left|\ |\alpha_{x}\!-\alpha_{y}|-2|\beta_{x}\!-\beta_{y}|\ \right|}$
    $\stackrel{{\scriptstyle\mathrm{(vi)}}}{{\leq}}\textstyle{|\alpha_{x}\!-\alpha_{y}|+(\frac{1}{2})^{n}|\alpha_{x}\!-\alpha_{y}|}$
[…]
    where the inequality labeled (vi) is derived using the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$, while the inequality labeled (vii) follows directly from (17).
To show (C1), from (16) and (17), we observe that in both the subcases
[…]
$\stackrel{{\scriptstyle\mathrm{(viii)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}}$
$\geq\textstyle{0.06-3\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where the inequality labeled (viii) is obtained using the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$.
For (C3), using Lemma 4(c),(e), we have
[…]
$=\textstyle{\left|(\beta_{x}\!-\gamma_{y})+(\frac{1}{2})^{n-1}(\beta_{x}\!-\gamma_{y})+(\frac{1}{2})^{n}(\alpha_{x}\!-\delta_{y})+n(\frac{1}{2})^{n}\right|}$
$\leq\textstyle{|\beta_{x}-\gamma_{y}|+(\frac{1}{2})^{n-1}|\beta_{x}-\gamma_{y}|+(\frac{1}{2})^{n}|\alpha_{x}\!-\delta_{y}|+n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(ix)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(x)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widehat{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
Using (10), (12), and the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$, we establish the inequality labeled (ix). Meanwhile, (18) leads to the inequality labeled (x).
B-B Proof of (C1)–(C3) in Case 3
For (C1), using Lemma 4(b), we have
$\quad\ |\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|$ $=\textstyle{|2(\alpha_{x}-\alpha_{y})+(\frac{1}{2})^{n}(\delta_{x}-\delta_{y})+\gamma_{x}-\gamma_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xi)}}}{{=}}\textstyle|2(\alpha_{x}-\alpha_{y})-(\frac{1}{2})^{n}(\alpha_{x}-\alpha_{y})-(\beta_{x}-\beta_{y})|$
$\geq\textstyle{\bigl(2-(\frac{1}{2})^{n}\bigr)|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\beta_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xii)}}}{{\geq}}\textstyle{\bigl(1-(\frac{1}{2})^{n}\bigr)|\alpha_{x}-\alpha_{y}|}$ (19)
$\stackrel{{\scriptstyle\mathrm{(xiii)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}-n(\frac{1}{2})^{n}}$
$\geq\textstyle{0.06-3\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where the equality labeled (xi) is derived from (13) and (14). The inequality labeled (xii) follows from the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$. Finally, the inequality labeled (xiii) is obtained using (9) and the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$.
For (C2), using Lemma 4(d), we have
$\quad\ |\beta_{\widecheck{x}}-\beta_{\widecheck{y}}|$ $=\textstyle{|\gamma_{x}\!-\gamma_{y}+(\frac{1}{2})^{n-1}(\beta_{x}\!-\beta_{y})+(\frac{1}{2})^{n}(\delta_{x}\!-\delta_{y})|}$
$\stackrel{{\scriptstyle\mathrm{(xiv)}}}{{=}}\textstyle{|\beta_{x}\!-\!\beta_{y}\!-(\frac{1}{2})^{n-1}(\beta_{x}\!-\!\beta_{y})\!+(\frac{1}{2})^{n}(\alpha_{x}\!-\!\alpha_{y})|}$
$\leq\textstyle{\left(1-(\frac{1}{2})^{n-1}\right)|\beta_{x}-\beta_{y}|+(\frac{1}{2})^{n}|\alpha_{x}-\alpha_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xv)}}}{{\leq}}\textstyle{\bigl(1-(\frac{1}{2})^{n}\bigr)|\alpha_{x}-\alpha_{y}|}$
$\stackrel{{\scriptstyle\mathrm{(xvi)}}}{{\leq}}|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|$
where the equality (xiv) is obtained directly from (13) and (14). The inequality labeled (xv) is established using the induction hypothesis that (C2) applies to $\mathrm{RM}(m,1)$, while inequality (xvi) results from (19).
For (C3), using Lemma 4(d),(f), we have
$|\beta_{\widecheck{x}}-\gamma_{\widecheck{y}}|$
$=\textstyle{|\gamma_{x}-\beta_{y}+(\frac{1}{2})^{n-1}(\beta_{x}-\gamma_{y})+(\frac{1}{2})^{n}(\delta_{x}-\alpha_{y})-n(\frac{1}{2})^{n}|}$
$\stackrel{{\scriptstyle\mathrm{(xvii)}}}{{\leq}}\textstyle{|\beta_{x}-\gamma_{y}|+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xviii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xix)}}}{{\leq}}\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|+n(\frac{1}{2})^{n}-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widecheck{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where the inequality (xvii) is obtained using (10), (12), and (15). The inequality labeled (xviii) is established by applying the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$, while inequality (xix) results from (9) and (19).
B-C Proof of (C1)–(C3) in Case 4
For (C1), using lemma 4(a),(b), we have
$|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|$
$=\textstyle{|2(\alpha_{x}-\alpha_{y})+\beta_{x}-\gamma_{y}+(\frac{1}{2})^{n}(\alpha_{x}-\delta_{y})+n(\frac{1}{2})^{n}|}$
$\geq\textstyle{2|\alpha_{x}-\alpha_{y}|-|\beta_{x}-\gamma_{y}|-(\frac{1}{2})^{n}|\alpha_{x}-\delta_{y}|-n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xx)}}}{{\geq}}\textstyle{|\alpha_{x}\!-\alpha_{y}|+0.06-4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}-2n(\frac{1}{2})^{n}}$ (20)
$\stackrel{{\scriptstyle\mathrm{(xxi)}}}{{\geq}}\textstyle{0.06-3\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+0.06-4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
$\stackrel{{\scriptstyle\mathrm{(xxii)}}}{{\geq}}0.12-0.007=0.113$
where the inequality labeled (xx) is derived using (10) along with the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$. The inequality labeled (xxi) is established based on the induction hypothesis that (C1) holds for $\mathrm{RM}(m,1)$. Finally, inequality (xxii) follows from the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq\sum\limits_{k=16}^{\infty}k\,(\frac{1}{2})^{k}=17\cdot(\frac{1}{2})^{15}<0.001$.
For (C2), using lemma 4(c),(d), we have
$|\beta_{\widehat{x}}-\beta_{\widecheck{y}}|$
$=\textstyle{|\beta_{x}\!-\gamma_{y}+(\frac{1}{2})^{n-1}(\beta_{x}\!-\beta_{y})+(\frac{1}{2})^{n}(\alpha_{x}\!-\delta_{y})+n(\frac{1}{2})^{n}|}$
$\stackrel{{\scriptstyle\mathrm{(xxiii)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+3n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xxiv)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+5n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.12+8\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
$\stackrel{{\scriptstyle\mathrm{(xxv)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|}$
where to obtain the inequality labeled (xxiii), we used (10), (11) and the induction hypothesis that (C3) holds for $\mathrm{RM}(m,1)$. The inequality labeled (xxiv) follows directly from (20), whereas we get the inequality labeled (xxv) based on the fact that $\sum\limits_{k=16}^{n}k\,(\frac{1}{2})^{k}\leq 0.001$.
For (C3), using lemma 4(c),(f), we have
$|\beta_{\widehat{x}}-\gamma_{\widecheck{y}}|$
$=\textstyle{|\beta_{x}-\beta_{y}+(\frac{1}{2})^{n-1}(\beta_{x}-\gamma_{y})+(\frac{1}{2})^{n}(\alpha_{x}-\alpha_{y})|}$
$\stackrel{{\scriptstyle\mathrm{(xxvi)}}}{{\leq}}\textstyle{|\alpha_{x}-\alpha_{y}|+2n(\frac{1}{2})^{n}}$
$\stackrel{{\scriptstyle\mathrm{(xxvii)}}}{{\leq}}\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n/2}k(\frac{1}{2})^{k}+2n(\frac{1}{2})^{n}+2n(\frac{1}{2})^{n}}$
$\leq\textstyle{|\alpha_{\widehat{x}}-\alpha_{\widecheck{y}}|-0.06+4\sum\limits_{k=16}^{n}k(\frac{1}{2})^{k}}$
where to obtain the inequality labeled (xxvi), we used (9), (12) and the induction hypothesis that (C2) holds for $\mathrm{RM}(m,1)$. Meanwhile, the inequality labeled (xxvii) follows directly from (20).